OpenStudy (anonymous):

The sum of two positive integers x and y is 90. Find the value of x that minimizes P=x^3-90xy. (Only an algebraic solution is accepted.)

OpenStudy (amoodarya):

x+y=90 so put y=90-x into function derive it f'=0 find roots find min

OpenStudy (anonymous):

Do I take the derivative of x^3-90xy then find the roots of that?

OpenStudy (amoodarya):

f(x,y)=x^3-90xy apply y=90-x so f(x)=x^3-90x(90-x) then find f'=0 roots

OpenStudy (amoodarya):

the roots of f'=0 must be -30 ,90 f=3x^2-180x+90^2